# Generic graded contractions of Lie algebras

### v1

@@ 2. $(ii^{\prime})$
Two graded contractions $\varepsilon$ and $\varepsilon^{\prime}$ are called equivalent, written $\varepsilon\sim\varepsilon^{\prime}$, if, for any $G$-graded Lie algebra $\mathfrak{L}$, there exists an equivalence of graded Lie algebras $\varphi\colon\mathfrak{L}^{\varepsilon}\to\mathfrak{L}^{\varepsilon^{\prime}}$, i.e., a Lie algebra isomorphism mapping each homogeneous component of $\mathfrak{L}^{\varepsilon}$ onto a homogeneous component of $\mathfrak{L}^{\varepsilon^{\prime}}$: for any $g\in G$ there is $h\in G$ such that $\varphi\left(\mathfrak{L}^{\varepsilon}_{g}\right)=\mathfrak{L}^{\varepsilon^{\prime}}_{h}$.
Definition 10.
Two contractions $\varepsilon$ and $\varepsilon^{\prime}$ are called strongly equivalent, written $\varepsilon\approx\varepsilon^{\prime}$, if, for any $G$-graded Lie algebra $\mathfrak{L}$, there exists an isomorphism of $G$-graded Lie algebras (for short, a graded isomorphism) $\varphi\colon\mathfrak{L}^{\varepsilon}\to\mathfrak{L}^{\varepsilon^{\prime}}$, i.e., an isomorphism of Lie algebras mapping each homogeneous component of $\mathfrak{L}^{\varepsilon}$ onto the homogeneous component of $\mathfrak{L}^{\varepsilon^{\prime}}$ of the same degree: $\varphi\big{(}\mathfrak{L}^{\varepsilon}_{g}\big{)}=\mathfrak{L}^{\varepsilon^{\prime}}_{g}$ for any $g\in G$.
It is clear that any two strongly equivalent contractions are equivalent, but not conversely, see [9, Example 2.5].
For arbitrary two graded contractions, $\varepsilon$ and $\varepsilon^{\prime}$, one can define their product $\varepsilon\varepsilon^{\prime}\colon G\times G\to\mathbb{F}$ to be simply the point-wise product of functions,
@@
The equivalence via normalization implies the strong equivalence, $\varepsilon\sim_{n}\varepsilon^{\prime}\Rightarrow\varepsilon\approx\varepsilon^{\prime}$. The graded isomorphism $\varphi\colon\mathfrak{L}^{\varepsilon}\to\mathfrak{L}^{\varepsilon^{\prime}}$ is given by
@@
for all $G$-graded Lie algebras $\mathfrak{L}$. The converse implication, $\varepsilon\approx\varepsilon^{\prime}\Rightarrow\varepsilon\sim_{n}\varepsilon^{\prime}$, is conjectured in [39, Conjecture 2.15]. While $\mathfrak{L}^{\varepsilon}$ may be isomorphic to $\mathfrak{L}^{\varepsilon^{\prime}}$, for specific $\mathfrak{L}$, without $\varepsilon$ and $\varepsilon^{\prime}$ being equivalent via normalization (see, for example, [9, Example 2.12]), we are not aware of counterexamples to this conjecture. Moreover, we show that a version of this conjecture holds when the graded isomorphism for all $\mathfrak{L}$ is replaced by its functorial counterpart, i.e., natural graded isomorphism. This is discussed in detail in Section 5.
@@ 3 Graded contractions via group cohomology
In [28, Section 3] Moody and Patera initiated the study of “generic” $G$-graded Lie algebra contractions, which they defined as functions $\varepsilon\colon G\times G\to\mathbb{F}$ satisfying conditions $(i^{\prime})$ and $(ii^{\prime})$ of Proposition 5, from the perspective of semigroup cohomology. Indeed, since the field $\mathbb{F}$ is a semigroup with respect to multiplication, it can be regarded as a $G$-semimodule with trivial action. In view of conditions $(i^{\prime})$ and $(ii^{\prime})$, a contraction $\varepsilon$ is a symmetric two-cocycle with values in this $G$-semimodule, moreover the equivalence via normalization coincides with the relation of “being cohomologous”. Since cohomology theory is more developed for groups than semigroups, we will follow a group-cohomology approach, which was inspired by the work of Weimar–Woods [39], even though group cohomology does not appear there explicitly.
@@
To find the set $\mathcal{S}(G)$, we develop an efficient approach based on some elementary facts from commutative algebra and affine algebraic geometry. It relies on the efficient algorithms of computing Gröbner bases and taking quotient rings in the computer algebra system Singular [7]. We discuss this algorithm in detail in Section 4.
@@
In [39], only supports of contractions, i.e. $S\in\mathcal{S}(G)$, were considered, but the definitions make sense for any subset of $\mathcal{P}_{G}$. Also, in [39], the relations $r_{1}=r_{2}$ of degrees $1$ and $2$ were excluded from higher-order identities because, for $S\in\mathcal{S}(G)$, they belong to $\langle\mathcal{R}_{S}\rangle$, i.e., follow from surviving defining relations. Indeed, a surviving identity of degree $1$ has the form $\{g_{1},h_{1}\}=\{g_{2},h_{2}\}$, where both sides are distinct and belong to $S$. Consider the group homomorphism $f\colon\mathcal{F}_{G}\to\mathbb{Z}G$ defined by $\{g,h\}\mapsto g+h-gh$. Since $f$ sends $\mathcal{R}_{G}$ to zero, it must kill all elements of $\langle\mathcal{R}_{G}\rangle$, in particular $\{g_{1},h_{1}\}\{g_{2},h_{2}\}^{-1}$, so we have $g_{1}+h_{1}-g_{1}h_{1}=g_{2}+h_{2}-g_{2}h_{2}$. This implies that the degree $1$ identity has the form $\{e,e\}=\{e,g\}$ with both sides belonging to $S$, so it follows from the surviving defining relation $\{e,e\}\{e,g\}=\{e,g\}\{e,g\}$. Similar considerations work for degree $2$ showing that the surviving degree $2$ identities are in fact surviving defining relations.
@@
For any abelian group $A$, viewed as a trivial $G$-module, the set of symmetric two-cocycles $Z^{2}_{sym}(G,A)$ with values in $A$ can be identified with ${\rm Hom}(\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle,A)$. It is well-known that
$Z^{2}_{sym}(G,A)/B^{2}(G,A)\simeq{\rm Ext}(G,A),$ (4)
where ${\rm Ext}(G,A)={\rm Ext}_{\mathbb{Z}}(G,A)$, which can be computed using the cochain complex ${\rm Hom}_{\mathbb{Z}}(C_{\bullet},A)$ for any free resolution $C_{\bullet}$ of the abelian group $G$.
@@
The exactness at the terms $C_{0}$ and $C_{1}$ is clear. To show exactness at $C_{2}$, we can apply the exact faithful functor ${\rm Hom}(\cdot,\mathbb{Q}/\mathbb{Z})$ and see that the resulting sequence is exact at ${\rm Hom}(C_{2},\mathbb{Q}/\mathbb{Z})$, i.e. $Z^{2}_{sym}(G,\mathbb{Q}/\mathbb{Z})=B^{2}(G,\mathbb{Q}/\mathbb{Z})$, because, in view of (4), the quotient $Z^{2}_{sym}(G,\mathbb{Q}/\mathbb{Z})/B^{2}(G,\mathbb{Q}/\mathbb{Z})$ is ${\rm Ext}(G,\mathbb{Q}/\mathbb{Z})$, which is trivial for any $G$. ∎
@@
It is claimed in [39, Theorem 6.5] that if $S$ belongs to $\mathcal{S}(G)$ then the group $(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle$ is also free (of rank $Q=N^{\prime\prime}-N^{\prime}$), but the proof appears to be incorrect (in the last paragraph of the proof, a function taking values in roots of unity is treated as a constant). Since we know that $\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})$ is free, the freeness of $(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle$ is equivalent to the freeness of $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$. We carried out numerical experiments for small groups $G$ and verified that the group $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ is free not only for $S\in\mathcal{S}(G)$, but for all subsets $S$ of $\mathcal{P}_{G}$.
@@
The group of equivalence classes of these contractions is isomorphic to the quotient by the image of $B^{2}(G,\mathbb{F}^{\times})$ under the restriction map ${\rm Fun}(\mathcal{P}_{G},\mathbb{F}^{\times})\to{\rm Fun}(S,\mathbb{F}^{\times})$, where the latter can be identified with the group ${\rm Hom}(\mathcal{F}_{S},\mathbb{F}^{\times})$ since $\mathcal{F}_{S}$ is the free abelian group generated by $S$. For any abelian group $A$, we denote by $B^{2}_{S}(A)$ the image of $B^{2}(G,A)$ in ${\rm Hom}(\mathcal{F}_{S},A)$, which lies in ${\rm Hom}(\mathcal{Z}_{S},A)$. We also denote
$H^{2}_{S}(A):={\rm Hom}(\mathcal{Z}_{S},A)/B^{2}_{S}(A).$
Then, for $S\in\mathcal{S}(G)$, the group $H^{2}_{S}(\mathbb{F}^{\times})={\rm Hom}(\mathcal{Z}_{S},\mathbb{F}^{\times})/B^{2}_{S}(\mathbb{F}^{\times})$ classifies contractions with support $S$ up to equivalence via normalization.
We want to compare the groups $\mathcal{Z}_{S}$ and $\mathcal{Z}_{G}:=\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$. There exists a unique group homomorphism $f\colon\mathcal{Z}_{S}\to\mathcal{Z}_{G}$ making the following diagram commutative:
00{0}${\langle\mathcal{R}_{S}\rangle}$${\mathcal{F}_{S}}$${\mathcal{Z}_{S}}$00{0}00{0}${\langle\mathcal{R}_{G}\rangle}$${\mathcal{F}_{G}}$${\mathcal{Z}_{G}}$00{0}$\scriptstyle{f}$
@@
$0\longrightarrow$ $\frac{{\rm Hom}(\mathcal{I}_{S},A)}{B^{2}_{S}(A)}\longrightarrow H^{2}_{S}(A)\longrightarrow{\rm Hom}(\mathcal{K}_{S},A)\longrightarrow 0,$ (5)
${\rm Ext}(G,A)\longrightarrow\frac{{\rm Hom}(\mathcal{I}_{S},A)}{B^{2}_{S}(A)}\longrightarrow{\rm Ext}(\mathcal{C}_{S},A)\longrightarrow 0,$ (6)
where the kernel of the map ${\rm Ext}(G,A)\to{\rm Hom}(\mathcal{I}_{S},A)/B^{2}_{S}(A)$ consists of the classes of symmetric two-cocycles whose restriction to $S$ is in $B^{2}_{S}(A)$. Moreover, (5) is split.
@@
Consider the short exact sequence $0\to\mathcal{K}_{S}\to\mathcal{Z}_{S}\;{\buildrel f\over{\to}}\;\mathcal{I}_{S}\to 0$. It splits since $\mathcal{I}_{S}$ is free, being a subgroup of $\mathcal{Z}_{G}$, which is a free abelian group by Corollary 17. This gives the following split exact sequence:
$0\to{\rm Hom}(\mathcal{I}_{S},A)\to{\rm Hom}(\mathcal{Z}_{S},A)\to{\rm Hom}(\mathcal{K}_{S},A)\to 0$ (7)
Since ${\rm Ext}(\mathcal{Z}_{G},A)=0$, the short exact sequence $0\to\mathcal{I}_{S}\to\mathcal{Z}_{G}\to\mathcal{C}_{S}\to 0$ gives
$0\to{\rm Hom}(\mathcal{C}_{S},A)\to{\rm Hom}(\mathcal{Z}_{G},A)\to{\rm Hom}(\mathcal{I}_{S},A)\to{\rm Ext}(\mathcal{C}_{S},A)\to 0$ (8)
The composition of ${\rm Hom}(\mathcal{Z}_{G},A)\to{\rm Hom}(\mathcal{I}_{S},A)$ from (8) and ${\rm Hom}(\mathcal{I}_{S},A)\to{\rm Hom}(\mathcal{Z}_{S},A)$ from (7) is the map $f^{*}$, which corresponds to the restriction of symmetric two-cocycles from $\mathcal{P}_{G}$ to $S$. By definition, it maps $B^{2}(G,A)$ onto $B^{2}_{S}(A)$. Recall that ${\rm Hom}(\mathcal{Z}_{G},A)/B^{2}(G,A)\simeq{\rm Ext}(G,A)$, so sequence (8) gives us (6) . Since $H^{2}_{S}(A)={\rm Hom}(\mathcal{Z}_{S},A)/B^{2}_{S}(A)$, sequence (7) gives us (5). ∎
Note that for contractions without zeros, Theorem 19 implies that $H^{2}_{S}(A)={\rm Ext}(G,A)$ as expected, since in this case $S=\mathcal{P}_{G}$ and therefore $f$ is the identity map.
@@
In the terminology of [39], the subgroup ${\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})/B^{2}_{S}(\mathbb{F}^{\times})$ in sequence (5) corresponds to the classes of contractions with support $S$ that satisfy all surviving higher-order identities, i.e., take value $1$ when applied to the elements of the subgroup $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ of $\mathcal{F}_{S}$.
The following corollary was obtained in the case $\mathbb{F}=\mathbb{C}$ in [39, Theorem 7.1].
@@
If $\mathbb{F}$ is algebraically closed then $H^{2}_{S}(\mathbb{F}^{\times})\simeq{\rm Hom}(\mathcal{K}_{S},\mathbb{F}^{\times})$.
@@
Since $\mathbb{F}^{\times}$ is a divisible group, both ${\rm Ext}(G,\mathbb{F}^{\times})$ and ${\rm Ext}(\mathcal{C}_{S},\mathbb{F}^{\times})$ are trivial, so sequence (6) implies ${\rm Hom}(\mathcal{I}_{S},A)/B^{2}_{S}(A)=0$. Consequently, sequence (5) gives the result. ∎
@@
If $\mathbb{F}$ is a real closed field (for example, $\mathbb{R}$), then the multiplicative group of $\mathbb{F}$ is the direct sum of {±$\{\pm 1\}\simeq\mathbb{Z}/2\mathbb{Z}$ and the divisible group of positive elements $\mathbb{F}_{>0}$, hence the group ${\rm Ext}(H,\mathbb{F}^{\times})$ is isomorphic to the dual group of $H_{[2]}$. Then sequence (6) takes the form
${\rm Hom}(G_{[2]},\{\pm 1\})\longrightarrow\frac{{\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})}{B^{2}_{S}(\mathbb{F}^{\times})}\longrightarrow{\rm Hom}((\mathcal{C}_{S})_{[2]},\{\pm 1\})\longrightarrow 0.$ (9)
@@
$H^{2}_{S}(\mathbb{F}^{\times})\simeq\frac{{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}{B^{2}_{S}(\{\pm 1\})}\oplus{\rm Hom}(\mathcal{K}_{S},\mathbb{F}^{\times}),$ (10)
where ${{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}/{B^{2}_{S}(\{\pm 1\})}$ is an elementary abelian $2$-group. If $G$ is finite, then the rank of this group (i.e., dimension as a vector space over the field of $2$ elements) is at most $\mathop{\rm rank}G_{[2]}+N^{\prime}$.
@@
Since $G_{[2]}$ and $(\mathcal{C}_{S})_{[2]}$ are elementary $2$-groups, (9) implies that ${\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})/B^{2}_{S}(\mathbb{F}^{\times})$ has period at most $4$. On the other hand, using the direct sum decomposition of $\mathbb{F}^{\times}$, we get that
$\frac{{\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})}{B^{2}_{S}(\mathbb{F}^{\times})}\simeq\frac{{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}{B^{2}_{S}(\{\pm 1\})}\oplus\frac{{\rm Hom}(\mathcal{I}_{S},\mathbb{F}_{>0})}{B^{2}_{S}(\mathbb{F}_{>0})}.$
@@
${\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})/B^{2}_{S}(\mathbb{F}^{\times})\simeq\frac{{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}{B^{2}_{S}(\{\pm 1\})}$

### v2

@@ 2. $(ii^{\prime})$
Two contractions $\varepsilon$ and $\varepsilon^{\prime}$ are called isomorphic, written $\varepsilon\simeq\varepsilon^{\prime}$, if, for any $G$-graded Lie algebra $\mathfrak{L}$, the contracted Lie algebras $\mathfrak{L}^{\varepsilon}$ and $\mathfrak{L}^{\varepsilon^{\prime}}$ are graded isomorphic.
For arbitrary graded contractions $\varepsilon$ and $\varepsilon^{\prime}$ one can define their product $\varepsilon\varepsilon^{\prime}\colon G\times G\to\mathbb{F}$ to be simply the point-wise product of functions,
@@
The equivalence via normalization implies isomorphism, $\varepsilon\sim_{n}\varepsilon^{\prime}\Rightarrow\varepsilon\simeq\varepsilon^{\prime}$. The graded isomorphism $\varphi\colon\mathfrak{L}^{\varepsilon}\to\mathfrak{L}^{\varepsilon^{\prime}}$ is given by
@@
for all $G$-graded Lie algebras $\mathfrak{L}$. The converse implication, $\varepsilon\simeq\varepsilon^{\prime}\Rightarrow\varepsilon\sim_{n}\varepsilon^{\prime}$, is conjectured in [48, Conjecture 2.15]. While $\mathfrak{L}^{\varepsilon}$ may be isomorphic to $\mathfrak{L}^{\varepsilon^{\prime}}$, for specific $\mathfrak{L}$, without $\varepsilon$ and $\varepsilon^{\prime}$ being equivalent via normalization (see, for example, [12, Example 2.12]), we are not aware of counterexamples to this conjecture. Moreover, we show that a version of this conjecture holds when the graded isomorphism for all $\mathfrak{L}$ is replaced by its functorial counterpart, i.e., natural graded isomorphism. This is discussed in detail in Section 5.
@@ 3 Graded contractions via group cohomology
In [36, Section 3] Moody and Patera initiated the study of “generic” $G$-graded Lie algebra contractions, which they defined as functions $\varepsilon\colon G\times G\to\mathbb{F}$ satisfying conditions $(i^{\prime})$ and $(ii^{\prime})$ of Proposition 5, from the perspective of semigroup cohomology. Indeed, since the field $\mathbb{F}$ is a semigroup with respect to multiplication, it can be regarded as a $G$-semimodule with trivial action. In view of conditions $(i^{\prime})$ and $(ii^{\prime})$, a contraction $\varepsilon$ is a symmetric two-cocycle with values in this $G$-semimodule, moreover the equivalence via normalization coincides with the relation of “being cohomologous”. Since cohomology theory is more developed for groups than semigroups, we will follow a group-cohomology approach, which was inspired by the work of Weimar–Woods [48], even though group cohomology does not appear there explicitly.
@@
To find the set $\mathcal{S}(G)$, we develop an approach based on some elementary facts from commutative algebra and affine algebraic geometry. It relies on the efficient algorithms of computing Gröbner bases and taking quotient rings in the computer algebra system Singular [10]. We discuss this algorithm in detail in Section 4.
@@
In [48], only supports of contractions, i.e., $S\in\mathcal{S}(G)$, were considered, but the definitions make sense for any subset of $\mathcal{P}_{G}$. Also, in [48], the relations $r_{1}=r_{2}$ of degrees $1$ and $2$ were excluded from higher-order identities because, for $S\in\mathcal{S}(G)$, they belong to $\langle\mathcal{R}_{S}\rangle$, i.e., follow from surviving defining relations. Indeed, a surviving identity of degree $1$ has the form $\{g_{1},h_{1}\}=\{g_{2},h_{2}\}$, where both sides are distinct and belong to $S$. Consider the group homomorphism $f\colon\mathcal{F}_{G}\to\mathbb{Z}G$ defined by $\{g,h\}\mapsto g+h-gh$. Since $f$ sends $\mathcal{R}_{G}$ to zero, it must kill all elements of $\langle\mathcal{R}_{G}\rangle$, in particular $\{g_{1},h_{1}\}\{g_{2},h_{2}\}^{-1}$, so we have $g_{1}+h_{1}-g_{1}h_{1}=g_{2}+h_{2}-g_{2}h_{2}$. This implies that the degree $1$ identity has the form $\{e,e\}=\{e,g\}$ with both sides belonging to $S$, so it follows from the surviving defining relation $\{e,e\}\{e,g\}=\{e,g\}\{e,g\}$. Similar considerations work for degree $2$ showing that the surviving degree $2$ identities are in fact surviving defining relations.
@@
For any abelian group $A$, viewed as a trivial $G$-module, the set of symmetric two-cocycles ${\rm Z}^{2}_{\rm sym}(G,A)$ with values in $A$ can be identified with ${\rm Hom}(\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle,A)$. It is well-known that
${\rm Z}^{2}_{\rm sym}(G,A)/{\rm B}^{2}(G,A)\simeq{\rm Ext}(G,A),$ (4)
see Section 2.1 for more details and references. The group ${\rm Ext}(G,A)$ can be computed using the cochain complex ${\rm Hom}(C_{\bullet},A)$ for any free resolution $C_{\bullet}$ of the abelian group $G$.
@@
The exactness at the terms $C_{0}$ and $C_{1}$ is clear. To show exactness at $C_{2}$, we can apply the exact faithful functor ${\rm Hom}(\cdot,\mathbb{Q}/\mathbb{Z})$, see, e.g., [45, Chapter 2, Exercise 2.5.1], and see that the resulting sequence is exact at ${\rm Hom}(C_{2},\mathbb{Q}/\mathbb{Z})$, i.e., ${\rm Z}^{2}_{\rm sym}(G,\mathbb{Q}/\mathbb{Z})={\rm B}^{2}(G,\mathbb{Q}/\mathbb{Z})$, because, in view of (4), the quotient ${\rm Z}^{2}_{\rm sym}(G,\mathbb{Q}/\mathbb{Z})/{\rm B}^{2}(G,\mathbb{Q}/\mathbb{Z})$ is ${\rm Ext}(G,\mathbb{Q}/\mathbb{Z})$, which is trivial for any $G$ since $\mathbb{Q}/\mathbb{Z}$ is a divisible group. ∎
@@
It is claimed in [48, Theorem 6.5] that if $S$ belongs to $\mathcal{S}(G)$ then the group $(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle$ is also free (of rank $Q=N^{\prime\prime}-N^{\prime}$), but the proof appears to be incorrect (in the last paragraph of the proof, a function taking values in roots of unity is treated as a constant). Since we know that $\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})$ is free, the freeness of $(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle$ is equivalent to the freeness of $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$. We carried out numerical experiments for small groups $G$ and verified that the group $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ is free not only for $S\in\mathcal{S}(G)$, but for all subsets $S$ of $\mathcal{P}_{G}$.
@@
The group of equivalence classes of these contractions is isomorphic to the quotient by the image of ${\rm B}^{2}(G,\mathbb{F}^{\times})$ under the restriction map ${\rm Fun}(\mathcal{P}_{G},\mathbb{F}^{\times})\to{\rm Fun}(S,\mathbb{F}^{\times})$, where ${\rm Fun}(X,Y)$ denotes the set of functions with domain $X$ and codomain $Y$, which becomes a group under the point-wise operation whenever $Y$ is a group. Since $\mathcal{F}_{S}$ is the free abelian group generated by $S$, ${\rm Fun}(S,\mathbb{F}^{\times})$ can be identified with the group ${\rm Hom}(\mathcal{F}_{S},\mathbb{F}^{\times})$. For any abelian group $A$, we denote by ${\rm B}^{2}_{S}(A)$ the image of ${\rm B}^{2}(G,A)$ in ${\rm Hom}(\mathcal{F}_{S},A)$, which lies in ${\rm Hom}(\mathcal{Z}_{S},A)$. We also denote
${\rm H}^{2}_{S}(A):={\rm Hom}(\mathcal{Z}_{S},A)/{\rm B}^{2}_{S}(A).$
Then, for $S\in\mathcal{S}(G)$, the group ${\rm H}^{2}_{S}(\mathbb{F}^{\times})={\rm Hom}(\mathcal{Z}_{S},\mathbb{F}^{\times})/{\rm B}^{2}_{S}(\mathbb{F}^{\times})$ classifies contractions with support $S$ up to equivalence via normalization.
We want to compare the groups $\mathcal{Z}_{S}$ and $\mathcal{Z}_{G}:=\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$. Since the embedding $\mathcal{F}_{S}\hookrightarrow\mathcal{F}_{G}$ maps $\langle\mathcal{R}_{S}\rangle$ into $\langle\mathcal{R}_{G}\rangle$, there exists a unique group homomorphism $f\colon\mathcal{Z}_{S}\to\mathcal{Z}_{G}$ making the following diagram commutative:
[…]
@@
$0\longrightarrow$ $\frac{{\rm Hom}(\mathcal{I}_{S},A)}{{\rm B}^{2}_{S}(A)}\longrightarrow{\rm H}^{2}_{S}(A)\longrightarrow{\rm Hom}(\mathcal{K}_{S},A)\longrightarrow 0,$ (5)
${\rm Ext}(G,A)\longrightarrow\frac{{\rm Hom}(\mathcal{I}_{S},A)}{{\rm B}^{2}_{S}(A)}\longrightarrow{\rm Ext}(\mathcal{C}_{S},A)\longrightarrow 0,$ (6)
where the kernel of the map ${\rm Ext}(G,A)\to{\rm Hom}(\mathcal{I}_{S},A)/{\rm B}^{2}_{S}(A)$ consists of the classes of symmetric two-cocycles whose restriction to $S$ is in ${\rm B}^{2}_{S}(A)$. Moreover, (5) is split.
@@
Consider the short exact sequence $0\to\mathcal{K}_{S}\to\mathcal{Z}_{S}\;{\mathrel{\mathop{\kern 0.0pt\to}\limits^{f}}}\;\mathcal{I}_{S}\to 0$. It splits since $\mathcal{I}_{S}$ is free, being a subgroup of $\mathcal{Z}_{G}$, which is a free abelian group by Corollary 17. This gives the following split exact sequence:
$0\to{\rm Hom}(\mathcal{I}_{S},A)\to{\rm Hom}(\mathcal{Z}_{S},A)\to{\rm Hom}(\mathcal{K}_{S},A)\to 0.$ (7)
The short exact sequence $0\to\mathcal{I}_{S}\to\mathcal{Z}_{G}\to\mathcal{C}_{S}\to 0$ induces the long exact sequence in cohomology, see, e.g., [33, Chapter II.4], which, given ${\rm Ext}(\mathcal{Z}_{G},A)=0$, takes the following form:
$0\to{\rm Hom}(\mathcal{C}_{S},A)\to{\rm Hom}(\mathcal{Z}_{G},A)\to{\rm Hom}(\mathcal{I}_{S},A)\to{\rm Ext}(\mathcal{C}_{S},A)\to 0.$ (8)
The composition of ${\rm Hom}(\mathcal{Z}_{G},A)\to{\rm Hom}(\mathcal{I}_{S},A)$ from (8) and ${\rm Hom}(\mathcal{I}_{S},A)\to{\rm Hom}(\mathcal{Z}_{S},A)$ from (7) is the map $f^{*}$, which corresponds to the restriction of symmetric two-cocycles from $\mathcal{P}_{G}$ to $S$. By definition, it maps ${\rm B}^{2}(G,A)$ onto ${\rm B}^{2}_{S}(A)$. Recall that ${\rm Hom}(\mathcal{Z}_{G},A)/{\rm B}^{2}(G,A)\simeq{\rm Ext}(G,A)$, so sequence (8) gives us (6) . Since ${\rm H}^{2}_{S}(A)={\rm Hom}(\mathcal{Z}_{S},A)/{\rm B}^{2}_{S}(A)$, sequence (7) gives us (5). ∎
Note that for contractions without zeros, Theorem 19 implies that ${\rm H}^{2}_{S}(A)={\rm Ext}(G,A)$ as expected, since in this case $S=\mathcal{P}_{G}$ and therefore $f$ is the identity map.
@@
In the terminology of [48], the subgroup ${\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})/{\rm B}^{2}_{S}(\mathbb{F}^{\times})$ in sequence (5) corresponds to the classes of contractions with support $S$ that satisfy all surviving higher-order identities, i.e., take value 11 when applied to the elements of the subgroup $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ of $\mathcal{F}_{S}$.
The following corollary was obtained in the case $\mathbb{F}=\mathbb{C}$ in [48, Theorem 7.1].
@@
If $\mathbb{F}$ is algebraically closed then ${\rm H}^{2}_{S}(\mathbb{F}^{\times})\simeq{\rm Hom}(\mathcal{K}_{S},\mathbb{F}^{\times})$.
@@
Since $\mathbb{F}^{\times}$ is a divisible group, both ${\rm Ext}(G,\mathbb{F}^{\times})$ and ${\rm Ext}(\mathcal{C}_{S},\mathbb{F}^{\times})$ are trivial, so sequence (6) implies ${\rm Hom}(\mathcal{I}_{S},A)/{\rm B}^{2}_{S}(A)=0$. Consequently, sequence (5) gives the result. ∎
@@
If $\mathbb{F}$ is a real closed field (for example, $\mathbb{R}$), then the multiplicative group of $\mathbb{F}$ is the direct sum of $\{\pm 1\}\simeq\mathbb{Z}/2\mathbb{Z}$ and the divisible group of positive elements $\mathbb{F}_{>0}$, hence the group ${\rm Ext}(H,\mathbb{F}^{\times})$ is isomorphic to the dual group of $H_{[2]}$. Then sequence (6) takes the form
${\rm Hom}(G_{[2]},\{\pm 1\})\longrightarrow\frac{{\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})}{{\rm B}^{2}_{S}(\mathbb{F}^{\times})}\longrightarrow{\rm Hom}((\mathcal{C}_{S})_{[2]},\{\pm 1\})\longrightarrow 0.$ (9)
@@
${\rm H}^{2}_{S}(\mathbb{F}^{\times})\simeq\frac{{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}{{\rm B}^{2}_{S}(\{\pm 1\})}\oplus{\rm Hom}(\mathcal{K}_{S},\mathbb{F}^{\times}),$ (10)
where ${{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}/{{\rm B}^{2}_{S}(\{\pm 1\})}$ is an elementary abelian $2$-group. If $G$ is finite, then the rank of this group (i.e., dimension as a vector space over the field of $2$ elements) is at most $\mathop{\rm rank}G_{[2]}+N^{\prime}$.
@@
Since $G_{[2]}$ and $(\mathcal{C}_{S})_{[2]}$ are elementary $2$-groups, (9) implies that ${\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})/{\rm B}^{2}_{S}(\mathbb{F}^{\times})$ has period at most $4$. On the other hand, using the direct sum decomposition of $\mathbb{F}^{\times}$, we get that
$\frac{{\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})}{{\rm B}^{2}_{S}(\mathbb{F}^{\times})}\simeq\frac{{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}{{\rm B}^{2}_{S}(\{\pm 1\})}\oplus\frac{{\rm Hom}(\mathcal{I}_{S},\mathbb{F}_{>0})}{{\rm B}^{2}_{S}(\mathbb{F}_{>0})}.$
@@
${\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})/{\rm B}^{2}_{S}(\mathbb{F}^{\times})\simeq\frac{{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}{{\rm B}^{2}_{S}(\{\pm 1\})}$